# Formatting Instructions For NeurIPS 2024

### v2

@@ Abstract
Large Language Models (LLMs) have recently demonstrated significant potential in the field of time series forecasting, offering impressive capabilities in handling complex temporal data. However, their robustness and reliability in real-world applications remain under-explored, particularly concerning their susceptibility to adversarial attacks. In this paper, we introduce a targeted adversarial attack framework for LLM-based time series forecasting. By employing both gradient-free and black-box optimization methods, we generate minimal yet highly effective perturbations that significantly degrade the forecasting accuracy across multiple datasets and LLM architectures. Our experiments, which include models like TimeGPT and LLM-Time with GPT-3.5, GPT-4, LLaMa, and Mistral, show that adversarial attacks lead to much more severe performance degradation than random noise, and demonstrate the broad effectiveness of our attacks across different LLMs. The results underscore the critical vulnerabilities of LLMs in time series forecasting, highlighting the need for robust defense mechanisms to ensure their reliable deployment in practical applications. The code is available at https://github.com/JohnsonJiang1996/AdvAttack_LLM4TS.
@@ 1 Introduction
Time series forecasting plays a pivotal role in numerous real-world applications, ranging from finance and healthcare to energy management and climate modeling. Accurately predicting temporal patterns in the data is crucial for informed decision-making in these domains liu2022universal . Recently, Large Language Models (LLMs), originally designed for Natural Language Processing (NLP) tasks, have demonstrated remarkable potential in handling time series forecasting challenges gruver2024large ; liu2024taming ; liu2024timecma ; tan2024language ; jin2023time . These models, including BERT devlin2018bert , GPT brown2020language ; achiam2023gpt , LLaMa touvron2023llama and their successors, leverage their powerful attention mechanisms and vast pre-training on diverse datasets to capture intricate temporal dependencies, making them highly effective for complex forecasting tasks.
@@
LLMs exhibit strong generalization capabilities across various types of time series data. Compared to traditional models like ARIMAkalpakis2001distance and Exponential Smoothing gardner1985exponential , as well as advanced deep learning models such as DNNs salinas2020deepar ; oreshkin2019n ; challu2023nhits , and Transformer-based architectureszhou2021informer ; wu2021autoformer ; liu2023itransformer ; zhou2022fedformer , LLMs excel in modeling long-term dependencies and capturing non-linear patterns within temporal sequences. This has resulted in impressive forecasting accuracy across applications ranging from energy consumption predictions to weather forecasting jin2023large ; jiang2024empowering .
@@
However, despite their success, the robustness and reliability of LLMs in real-world forecasting remain concerns, particularly their vulnerability to adversarial attacks is under-explored. Adversarial attacks introduce subtle, often imperceptible perturbations to input data, leading to significant and misleading changes in model predictions. While the susceptibility of machine learning models to such attacks has been well-explored in image processing and NLP domains xu2020adversarial ; morris2020textattack ; wei2018transferable , there is a noticeable gap in research on their impact on LLMs used for time series forecasting.
@@
Figure 1: Adversarial Black-box Attack for LLMs in Time Series Forecasting
@@
While adversarial attacks and defenses for deep neural networks have been extensively studied across various domains madry2017towards , executing adversarial attacks against LLMs in time series forecasting presents two significant challenges. First, to prevent information leakage, we cannot use ground truth values (i.e., future time steps) when attacking forecasting models. Second, LLMs must be treated as strict black-box systems due to the difficulty of accessing their internal workings and parameters.
@@
In this paper, we address this gap by proposing a gradient-free black-box attack that transforms the output of LLM-based forecasting models into a random walk, while investigating the vulnerabilities of large language models in time series forecasting. As depicted in Figure 1, we demonstrate that even minimal attack perturbations can cause substantial deviations in LLMs’ predictions. We evaluate two forms of LLM applications for time series forecasting, encompassing five sub-models, across five datasets from various real-world domains. Our findings reveal that LLMs, despite their advanced architectures, are indeed susceptible to adversarial manipulations in time series domain, leading to unstable and inaccurate forecasts. This underscores the urgent need to develop more robust LLMs that can withstand such attacks, ensuring their reliability in real-world applications.
@@ 2.1 Adversarial Attacks in Time Series Forecasting
Adversarial attacks in time series forecasting have emerged as a crucial area of research, exposing vulnerabilities in forecasting models. Unlike adversarial studies in static domains, such as object recognition or time series classification, adversarial attacks on time series forecasting cannot leverage ground truth data for perturbation generation due to the risk of information leakage liu2022practical . To address this challenge, surrogate techniques have been adopted liu2021spatially , which bypass the need for labels, as is done in traditional adversarial attack methods like the Fast Gradient Sign Method goodfellow2014explaining . Several studies have treated forecasting models as white-box systems to investigate the effects of adversarial attacks on commonly used models in time series forecasting, such as ARIMA, LSTMs, and Transformer-based models liu2022robust ; liu2023adversarial . These studies demonstrate that even small perturbations can severely impact these models, resulting in inaccurate forecasts. However, evaluating the vulnerability of LLM-based forecasting presents a significant challenge, as internal access is typically restricted, requiring these models to be treated as black-box systems.
@@ 2.2 Adversarial Attacks on LLMs
  • Jailbreak Prompting yu2024don ; wei2024jailbroken : Crafted prompts that bypass LLM guardrails, inducing unintended or harmful outputs by exploiting unconventional phrasing.
  • Prompt Injection greshake2023not ; xue2024trojllm ; liu2024automatic : Adversarial instructions embedded into benign prompts to manipulate LLM responses, highlighting their vulnerability to prompt manipulation.
  • White-box Gradient Attacks madry2017towards ; guo2021gradient : Using internal model parameters, attackers apply gradient-based methods to perturb inputs, significantly altering outputs with minimal changes.
  • Black-box Attacks sitawarin2024pal : Query-based attacks without model access, using techniques like Zeroth-Order Optimization to craft adversarial examples by estimating gradients.
  • Embedding Perturbations schwinn2024soft ; singh2024robustness : Subtle changes to input embeddings disrupt LLM’s internal representations, leading to erroneous outputs with minimal visible input alterations.
@@ 3.1 Formulations of LLM-based Time Series Forecasting
LLMs have shown promising performance in time series forecasting by leveraging their ability to perform next-token prediction, a technique originally developed for text-based tasks gruver2024large ; jin2023time . A typical LLM-based time series forecasting model, denoted as $f(\cdot)$, consists of two primary components: an embedding or tokenization module that encodes the time series data into a sequence of tokens, and a pre-trained LLM that autoregressively predicts the subsequent tokens. The embedding module translates the raw time series into a format suitable for the LLM, while the LLM captures the temporal dependencies and generates predictions based on its learned representations.
@@ 3.2 Threat model
In practical applications, it is infeasible to access the full set of detailed parameters of an LLM, leading the attacker to treat the target model as a black box. Additionally, obtaining the entire training dataset is impractical, meaning the attacker does not have access to this data. The attacker’s capabilities are summarized as follows: (i) no access to the training data, (ii) no access to internal information of the LLM-based forecasting model, and (iii) be able to query the target model.
4 Target Attack with Directional Gradient Approximation
@@
$\bm{g}_{t-T+1:t}=\frac{\mathcal{L}\left(\mathcal{Y}-f\left(\mathbf{X}_{t-T+1:t}+\bm{\theta}_{t-T+1:t}\right)\right)-\mathcal{L}\left(\mathcal{Y}-f\left(\mathbf{X}_{t-T+1:t}\right)\right)}{\bm{\theta}_{t-T+1:t}}.$ (5)
@@
$\mathbf{X^{\prime}}_{t-T+1:t}=\mathbf{X}_{t-T+1:t}+\bm{\rho}_{t-T+1:t}=\mathbf{X}_{t-T+1:t}+\epsilon\cdot\text{sign}\left(\bm{g}_{t-T+1:t}\right),$ (6)
@@ 5.1 Datasets
  • ETTh1 and ETTh2 (Electricity Transformer Temperature Hourly) zhou2021informer : These datasets consist of two years of hourly recorded data from electricity transformers, capturing temperature and power consumption variables.
  • IstanbulTrafficgruver2024large : This dataset contains hourly measurements of road traffic volumes across different sensors. It captures temporal dependencies related to traffic patterns, making it ideal for testing models on dynamic and fluctuating time series data.
  • Weather zhou2021informer : This dataset comprises meteorological data, including variables such as temperature, humidity, and wind speed, recorded hourly. It provides a challenging forecasting task due to the inherent variability and complexity of weather patterns.
  • Exchange lai2018modeling : This dataset consists of daily exchange rates from eight foreign countries—Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore—covering the period from 1990 to 2016.
These diverse datasets allow us to evaluate the robustness of LLMs across different types of temporal dynamics and forecasting challenges. In our experiments, 50% of the data is used for training, while the remaining data is split evenly: 25% for validation and 25% for testing. It should be noted that the attacker does not access either the training or validation part. We use a 96-step historical time window as input to the forecasting model, which predicts the subsequent 48-step future values.
@@ 5.2 Target Models
To assess the impact of adversarial attacks on LLMs for time series forecasting, we selected two state-of-the-art LLM-based forecasting models as baselines, which together represent two common forms of LLM application for time series tasks:
  • TimeGPT garza2023timegpt : A large model specifically pre-trained with a vast amount of time series data. TimeGPT uses advanced attention mechanisms and temporal encoding to capture complex patterns in sequential data, making it a leading LLM designed explicitly for time series forecasting. Its pre-training, which is conducted from scratch using vast amounts of time series data, allows it to serve as a robust and versatile tool for a wide range of time-dependent applications.
  • LLMTime gruver2024large : This model treats time series forecasting as a next-token prediction task, using LLM architectures like GPT and LLaMa. By converting time series data into numerical sequences, LLM-Time enables these models to apply their sequence prediction strengths to time series. To test the robustness of our adversarial attacks, we experimented with base models including GPT-3.5, GPT-4, LLaMa, and Mistral, assessing their resilience when adapted from natural language processing to time series forecasting.
  • TimeLLM jin2023time : TimeLLM presents a novel approach for time series forecasting by adapting LLMs with reprogramming input time series data into textual representations that are more compatible with LLMs, allowing the models to perform time series forecasting tasks without altering their pre-trained structures. The key innovation is the Prompt-as-Prefix (PaP) technique, which augments input context to guide the LLM in transforming reprogrammed data into accurate forecasts.
@@ 5.4 Overall Comparison
As shown in Table 1, the experimental results demonstrate that the designed adversarial attacks significantly degraded forecasting performance across all datasets, as indicated by increased MSE and MAE values. Compared to GWN of the same perturbation intensity, our attacks had a much more detrimental effect on the models’ predictions.
For TimeGPT, which is pre-trained with large-scale time series data, the adversarial attack led to a sharp rise in forecasting errors, demonstrating that even models specifically built for time series forecasting are vulnerable. For LLM-Time, which includes GPT-3.5, GPT-4, LLaMa, and Mistral as base models, the adversarial attack was even more pronounced. As illustrated in Figure LABEL:fig:lines, the attack caused a clear divergence between the forecasted values and the true time series, with all different variants of LLM-Time exhibiting larger deviations compared to GWN. GPT-3.5 and GPT-4, in particular, showed significant susceptibility, with their errors increasing substantially under adversarial conditions.
@@ 6 Conclusion
In this study, we demonstrated the significant vulnerabilities of LLM-based models for time series forecasting to adversarial attacks. Through a comprehensive evaluation of TimeGPT and LLM-Time (with GPT-3.5, GPT-4, LLaMa, and Mistral as base models), we found that targeted adversarial perturbations, generated using Directional Gradient Approximation (DGA), caused substantial increases in prediction errors. These attacks were far more damaging than Gaussian White Noise (GWN) of similar intensity, highlighting the precision and effectiveness of the adversarial strategy.
The experimental results revealed that both large, pre-trained models like TimeGPT and fine-tuned models such as LLM-Time are highly susceptible to adversarial manipulation. The proposed attack can significantly degrade model performance across various datasets. This poses serious challenges for the deployment of LLMs in real-world time series applications, where reliability is critical.
Our findings emphasize the need for future research to focus on developing robust defense mechanisms to mitigate adversarial threats and enhance the resilience of LLM-based time series forecasting models. Without such protections, these models remain vulnerable to attacks that could undermine their practical utility in high-stakes environments. In addition, future studies should compare vulnerabilities of LLM with lighter models.

### v3

# Adversarial Vulnerabilities in Large Language Models for Time Series Forecasting

@@ Abstract
Large Language Models (LLMs) have recently demonstrated significant potential in time series forecasting, offering impressive capabilities in handling complex temporal data. However, their robustness and reliability in real-world applications remain under-explored, particularly concerning their susceptibility to adversarial attacks. In this paper, we introduce a targeted adversarial attack framework for LLM-based time series forecasting. By employing both gradient-free and black-box optimization methods, we generate minimal yet highly effective perturbations that significantly degrade the forecasting accuracy across multiple datasets and LLM architectures. Our experiments, which include models like LLMTime with GPT-3.5, GPT-4, LLaMa, and Mistral, TimeGPT, and TimeLLM show that adversarial attacks lead to much more severe performance degradation than random noise, and demonstrate the broad effectiveness of our attacks across different LLMs. The results underscore the critical vulnerabilities of LLMs in time series forecasting, highlighting the need for robust defense mechanisms to ensure their reliable deployment in practical applications.
@@ 1 Introduction
Time series forecasting plays a pivotal role in numerous real-world applications, ranging from finance and healthcare to energy management and climate modeling. Accurately predicting temporal patterns in the data is crucial for informed decision-making in these domains [1]. Recently, Large Language Models (LLMs), originally designed for Natural Language Processing (NLP) tasks, have demonstrated remarkable potential in handling time series forecasting challenges [2, 3, 4, 5, 6]. These models, including BERT [7], GPT [8, 9], LLaMa [10] and their successors, leverage their powerful attention mechanisms and vast pre-training on diverse datasets to capture intricate temporal dependencies, making them highly effective for complex forecasting tasks.
@@
LLMs exhibit strong generalization capabilities across various types of time series data. Compared to traditional models like ARIMA[11] and Exponential Smoothing [12], as well as advanced deep learning models such as DNNs [13, 14, 15], and Transformer-based architectures[16, 17, 18, 19], LLMs excel in modeling long-term dependencies and capturing non-linear patterns within temporal sequences. This has resulted in impressive forecasting accuracy across applications ranging from energy consumption predictions to weather forecasting [20, 21].
@@
However, despite their success, the robustness and reliability of LLMs in real-world forecasting remain concerns, particularly their vulnerability to adversarial attacks is under-explored. Adversarial attacks introduce subtle, often imperceptible perturbations to input data, leading to significant and misleading changes in model predictions. While the susceptibility of machine learning models to such attacks has been well-explored in image processing and NLP domains [22, 23, 24], there is a noticeable gap in research on their impact on LLMs used for time series forecasting.
@@
Figure 1: Adversarial Black-box Attack for LLMs in Time Series Forecasting.
@@
While adversarial attacks and defenses for deep neural networks have been extensively studied across various domains [25], executing adversarial attacks against LLMs in time series forecasting presents two significant challenges. First, to prevent information leakage, we cannot use ground truth values (i.e., future time steps) when attacking forecasting models. Second, LLMs must be treated as strict black-box systems due to the difficulty of accessing their internal workings and parameters.
@@
In this paper, we address this gap by proposing a gradient-free black-box attack that transforms the output of LLM-based forecasting models into a random walk, while investigating the vulnerabilities of large language models in time series forecasting. As depicted in Figure 1, we demonstrate that even minimal attack perturbations can cause substantial deviations in LLMs’ predictions. We evaluate three forms of LLM applications for time series forecasting, encompassing six sub-models as well as two non-LLM models, across five datasets from various real-world domains. Our findings reveal that LLMs, despite their advanced architectures, are indeed susceptible to adversarial manipulations in time series domain, leading to unstable and inaccurate forecasts. This underscores the urgent need to develop more robust LLMs that can withstand such attacks, ensuring their reliability in real-world applications.
@@ 2.1 Adversarial Attacks in Time Series Forecasting
Adversarial attacks in time series forecasting have emerged as a crucial area of research, exposing vulnerabilities in forecasting models. Unlike adversarial studies in static domains, such as object recognition or time series classification, adversarial attacks on time series forecasting cannot leverage ground truth data for perturbation generation due to the risk of information leakage [26]. To address this challenge, surrogate techniques have been adopted [27], which bypass the need for labels, as is done in traditional adversarial attack methods like the Fast Gradient Sign Method [28]. Several studies have treated forecasting models as white-box systems to investigate the effects of adversarial attacks on commonly used models in time series forecasting, such as ARIMA, LSTMs, and Transformer-based models [29, 30]. These studies demonstrate that even small perturbations can severely impact these models, resulting in inaccurate forecasts. However, evaluating the vulnerability of LLM-based forecasting presents a significant challenge, as internal access is typically restricted, requiring these models to be treated as black-box systems.
@@ 2.2 Adversarial Attacks on LLMs
  • Jailbreak Prompting [31, 32]: Crafted prompts that bypass LLM guardrails, inducing unintended or harmful outputs by exploiting unconventional phrasing.
  • Prompt Injection [33, 34, 35]: Adversarial instructions embedded into benign prompts to manipulate LLM responses, highlighting their vulnerability to prompt manipulation.
  • Gradient Attacks [25, 36]: Using internal model parameters, attackers apply gradient-based methods to perturb inputs, significantly altering outputs with minimal changes.
  • Embedding Perturbations [37, 38]: Subtle changes to input embeddings disrupt LLM’s internal representations, leading to erroneous outputs with minimal visible input alterations.
@@ 3.1 Formulations of LLM-based Time Series Forecasting
LLMs have shown promising performance in time series forecasting by leveraging their ability to perform next-token prediction, a technique originally developed for text-based tasks [2, 6]. A typical LLM-based time series forecasting model, denoted as $f(\cdot)$, consists of two primary components: an embedding or tokenization module that encodes the time series data into a sequence of tokens, and a pre-trained LLM that autoregressively predicts the subsequent tokens. The embedding module translates the raw time series into a format suitable for the LLM, while the LLM captures the temporal dependencies and generates predictions based on its learned representations.
@@ 3.2 Threat model
In practical applications, accessing the full set of detailed parameters of an LLM is typically infeasible, which forces the attacker to treat the target model as a black-box system. Additionally, acquiring the complete training dataset is impractical, meaning the attacker lacks access to this information as well. The attacker’s capabilities can be summarized as follows: (i) no access to the training data, (ii) no access to internal information of the LLM-based forecasting model, (iii) no access to ground truth, and (iv) the ability to query the target model.
The threat model for adversarial attacks against LLM-based time series forecasting underscores the complexity of this task. Unlike attacks on LLMs in static applications, the attacker here cannot leverage labels for crafting attacks. Furthermore, compared to attacks on non-LLM forecasting models, the internal details of LLMs are strictly inaccessible, prohibiting the use of white-box attack techniques. This restriction highlights the increased challenge of developing effective adversarial attacks in this context.
4 Directional Gradient Approximation
@@
$\begin{split}&\bm{g}_{t-T+1:t}=\\ &\frac{\mathcal{L}\left(\mathcal{Y}-f\left(\mathbf{X}_{t-T+1:t}+\bm{\theta}_{t-T+1:t}\right)\right)-\mathcal{L}\left(\mathcal{Y}-f\left(\mathbf{X}_{t-T+1:t}\right)\right)}{\bm{\theta}_{t-T+1:t}}.\end{split}$ (5)
@@
$\begin{split}\mathbf{X^{\prime}}_{t-T+1:t}&=\mathbf{X}_{t-T+1:t}+\bm{\rho}_{t-T+1:t}\\ &=\mathbf{X}_{t-T+1:t}+\epsilon\cdot\text{sign}\left(\bm{g}_{t-T+1:t}\right),\end{split}$ (6)
@@ 5.1 Datasets
  • ETTh1 and ETTh2 (Electricity Transformer Temperature Hourly) [16]: These datasets consist of two years of hourly recorded data from electricity transformers, capturing temperature and power consumption variables.
  • IstanbulTraffic[2]: This dataset contains hourly measurements of road traffic volumes across different sensors. It captures temporal dependencies related to traffic patterns, which are dynamic and fluctuating time series.
  • Weather [39]: This dataset comprises meteorological data, including variables such as temperature, humidity, and wind speed, recorded hourly. It provides a challenging forecasting task due to the inherent variability and complexity of weather patterns.
  • Exchange [40]: This dataset consists of daily exchange rates from eight foreign countries—Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore—covering the period from 1990 to 2016.
These diverse datasets allow us to evaluate the adversarial robustness of LLMs across different types of temporal dynamics and forecasting challenges. In our experiments, 50% of the data is used for training, while the remaining data is split evenly: 25% for validation and 25% for testing. We use a 96-step historical time window as input to the forecasting model, which predicts the subsequent 48-step future values. It should be noted that the attacker does not access either the training or validation part.
@@ 5.2 Target Models
To assess the impact of adversarial attacks on LLMs for time series forecasting, we selected three state-of-the-art LLM-based forecasting models as baselines, which together represent three common forms of LLM application for time series tasks:
  • TimeGPT [41]: A large model specifically pre-trained with a vast amount of time series data. It uses advanced attention mechanisms and temporal encoding to capture complex patterns in sequential data, making it a leading LLM designed explicitly for time series forecasting. Its pre-training, which is conducted from scratch using vast amounts of time series data, allows it to serve as a robust and versatile tool for a wide range of time-dependent applications.
  • LLMTime [2]: This model treats time series forecasting as a next-token prediction task, using LLM architectures like GPT and LLaMa. By converting time series data into numerical sequences, LLMTime enables these models to apply their sequence prediction strengths to time series. To test the robustness of our adversarial attacks, we experimented with base models including GPT-3.5, GPT-4, LLaMa, and Mistral, assessing their resilience when adapted from natural language processing to time series forecasting.
  • TimeLLM [6]: This model presents a novel approach for time series forecasting by adapting LLMs by reprogramming input time series into textual representations that are more compatible with LLMs, allowing the models to perform time series forecasting tasks without altering their pre-trained structures. The key innovation is the Prompt-as-Prefix (PaP) technique, which augments input context to guide the LLM in transforming reprogrammed data into accurate forecasts.
  • TimeNet[39] and iTransformer[18]are two supervised forecasting models that leverage the attention-based architecture of transformers, enabling them to capture long-term dependencies in time series data effectively. These models provide strong performance in time series forecasting tasks, and in this study, they are used in the performance and robustness comparison between LLM-based forecasting models with non-LLM models.
Both TimeGPT and LLMTime operate as zero-shot forecasters, making predictions for each time series without prior exposure to the dataset. In contrast, TimeLLM is fine-tuned using 10% of each dataset for every task. On the other hand, the two non-LLM forecasters, TimeNet and iTransformer, are trained using the full available training data for each dataset.
@@ 5.4 Overall Comparison
As shown in Table 1, the experimental results demonstrate that the designed adversarial attacks significantly degraded forecasting performance across all datasets, as indicated by increased MSE and MAE values. Compared to GWN of the same perturbation intensity, our attacks had a much more detrimental effect on the models’ predictions. For TimeGPT, which is pre-trained with large-scale time series data, the adversarial attack led to a sharp rise in forecasting errors, demonstrating that even models specifically built for time series forecasting are vulnerable. For LLMTime, which includes GPT-3.5, GPT-4, LLaMa, and Mistral as base models, the adversarial attack was even more pronounced.
Figure LABEL:fig:robustness_comparison demonstrates the robustness comparison between LLM-based forecasting models (LLMTime with GPT-4 and TimeGPT) and non-LLM models (iTransformer and TimeNet) under the proposed black-box adversarial attack, DGA. The larger blue areas in the radar charts for the LLM-based models indicate that they experience significantly higher increases in errors, across all datasets (ETTh1, ETTh2, Exchange, Traffic, and Weather). In contrast, the non-LLM models, iTransformer and TimeNet, exhibit much smaller error increases, suggesting that they are more robust to adversarial attacks. This analysis highlights that LLM-based models are generally less resilient than non-LLM models, making them more vulnerable to adversarial manipulations in time series forecasting.
As illustrated in Figure LABEL:fig:lines, the attack caused a clear divergence between the forecasted values and the true time series, with all different variants of LLMTime exhibiting larger deviations compared to GWN. LLMTime with GPT-3.5 in particular, showed significant susceptibility, with their errors increasing substantially under adversarial conditions.
@@ 5.6 Hyperparameter Study
5.7 Discussion of Mitigation Methods
Challenges of Adversarial Training in LLM4TS: Adversarial training effectively mitigates attacks but poses challenges for LLM-based time series forecasting due to the high computational costs of pretraining on large datasets. Its iterative nature further escalates costs, as it requires generating adversarial examples and optimizing against them during training. Retrofitting pre-trained models like TimeGPT with adversarial defenses is similarly impractical, often demanding months of computation and significant pipeline modifications.
Alternative Mitigation Strategies for LLM4TS: To overcome these challenges, preprocessing-based methods provide a practical alternative. Filter-based defenses reform time series data before forecasting, while machine learning-based anomaly detection identifies and filters adversarial inputs. These computationally efficient approaches are particularly effective against black-box adversarial attacks, offering a promising defense for LLM4TS models. Future work will focus on refining and evaluating these strategies.
@@ 6 Conclusion
In this study, we demonstrated the significant vulnerabilities of LLM-based models for time series forecasting to adversarial attacks. Through a comprehensive evaluation of TimeGPT and LLMTime (with GPT-3.5, GPT-4, LLaMa, and Mistral as base models), we found that targeted adversarial perturbations, generated using Directional Gradient Approximation (DGA), caused substantial increases in prediction errors. These attacks were far more damaging than Gaussian White Noise (GWN) of similar intensity, highlighting the precision and effectiveness of the adversarial strategy.
The experimental results revealed that both large, pre-trained models like TimeGPT and fine-tuned models such as LLMTime are highly susceptible to adversarial manipulation. The proposed attack can significantly degrade model performance across various datasets. This poses serious challenges for the deployment of LLMs in real-world time series applications, where reliability is critical.
Our findings emphasize the need for future research to focus on developing robust defense mechanisms to mitigate adversarial threats and enhance the resilience of LLM-based time series forecasting models. Without such protections, these models remain vulnerable to attacks that could undermine their practical utility in high-stakes environments.